factorize
(x-y+z)^2+2(x-y+z)(y-z+x)+(y-z-x)
1
Expert's answer
2012-06-19T10:53:09-0400
(x-y+z)^2+2(x-y+z)(y-z+x)+(y-z-x)=(x-y+z)^2+2(x-y+z)(y-z+x)-(-y+z+x)=(x-y+z)*(x-y+z+2*(y-z+x)-1)=(x-y+z)*(x-y+z+2y-2z+2x-1)=(x-y+z)(3x+y-z-1)